In 8 years Eli will be five times as old as he was 12 years ago. How old is he now?
Answer by harpazo(655)

Eli's current age = x
x + 8 = 5(x-12)

x + 8 = 5x -60
x-5x = -60-8
-4x = -68
x = -68/-4
x = 17
Eli is 17 years old.